# Quality-aware Masked Diffusion Transformer for Enhanced Music Generation

### v1

@@ 5 Experimental setup
5.1 Datasets
Train datasets. We used the following databases for our training: AudioSet Music Subset (ASM) (Gemmeke et al., 2017), MagnaTagTune (MTT) (Law et al., 2009), Million Song Dataset (MSD) (Bertin-Mahieux et al., 2011), Free Music Archive (FMA) (Defferrard et al., 2016), and a private dataset. Each track in these databases was clipped to 10-second segments and sampled at 16kHz to ensure uniformity across the dataset. The final training set was developed through a process of caption refinement, as detailed in Section 4.3. Finally, we got our training set totaling 12.5k hours of diverse music data. The specific composition of these datasets is further elaborated in the Appendix A.1.
Evaluation datasets. For comparison with prior work, we evaluate our model on the widely used MusicCaps benchmark (Agostinelli et al., 2023). MusicCaps consists of 5.5K $10.24$-second clips that include high-quality music descriptions from ten musicians.
5.2 Models and hyperparameters
@@
Each 10.24-second audio clip, sampled at 16 kHz, is initially transformed into a $64\times 1024$ mel-spectrogram with mel-bins of 64, hop-length of 160 and window length of 1024. Subsequently, this spectrogram is compressed into a $16\times 128$ latent representation $\mathcal{M}_{spec}$ using a Variational Autoencoder (VAE) pretrained with AudioLDM 2 (Liu et al., 2023a). This latent representation is utilized for adding noise and training purposes. Finally, we employ a pretrained Hifi-GAN (Kong et al., 2020) to reconstruct the waveform from the generated mel-spectrogram.
@@
Caption processing and conditioning. We utilize the LP-MusicCaps (Doh et al., 2023) caption model for ASM, FMA, and subsets of MTT and MSD that have weak or no captions. We use the LAION-CLAP (Wu* et al., 2023) checkpoint music_speech_audioset_epoch_15_esc_89.98.pt for text-to-text and text-to-audio similarity calculations. Thresholds are set at $\rho_{1}=\rho_{2}=0.1$ to ensure any generated text or original caption not aligned well with the corresponding waveform is filtered out. Additionally, after filtering, generated text that fall below a threshold of $\rho_{3}=0.25$ are merged with original tags with the prompt: Merge this music caption “generated caption” with the ground truth tags “original tags”. Do not add any imaginary elements and try to keep the modified caption refined and accurate. We use FLAN-T5-large (Peebles and Xie, 2023) as text encoder for all models.
@@
Diffusion backbone. We train our diffusion model with three backbones for comparison: U-Net (Ronneberger et al., 2015) based at 1.0B parameters, code based on AudioLDM (Liu et al., 2023c); and our proposed Quality-aware Masked Diffusion Transformer based at 675M parameters with $N=20$ encoder layers and $M=8$ decoder layers. Flash attention (Dao et al., 2022) from the xFormers package (Lefaudeux et al., 2022) is utilized to improve training speed and memory usage for self-attention and cross-attention stages. We study the impact of the patch size and overlap size in the Appendix A.2 , and apply a patch size of $1\times 4$ without overlap for the training of our final model. We train on 10-second audio crops sampled at random from the full track, maintaining a total batch size of 64 on four NVIDIA A100 80GB GPUs, with gradient clipping of 1.0, learning rate of 8e-5, 2000 steps of linear warmup without weight decay, and a condition drop of 0.1 during training. The final model was trained for a total of seven days. During inference, we use Denoising Diffusion Implicit Models (DDIM)  (Song et al., 2020) with 200 steps and a guidance scale of 3.5, consistent with AudioLDM  (Liu et al., 2023c).
5.3 Evaluation metrics
We evaluate the proposed method using objective metrics, including the Fréchet Audio Distance (FAD) (Kilgour et al., 2018), Kullback-Leibler Divergence (KL), Inception Score (IS). All above metrics are computed using the audioldm_eval library (Liu et al., 2023c), ensuring standardized and reliable evaluation. We also utilize pseudo-MOS scoring model (Ragano et al., 2023) to estimate generation quality, with more accurate assessments derived from subjective metrics.
@@ 5.3 Evaluation metrics
For the human studies, we ask human raters to evaluate two aspects of the audio samples (i) overall quality (Ovl), and (ii) relevance to the text input (Rel). For the overall quality test, raters were asked to rate the perceptual quality of the provided samples with 1 to 5. For the text relevance test, raters were asked to rate the match between audio and text on a scale from 1 to 5. Our evaluators consist of people from various backgrounds, including professional music producers, video editors who need to use soundtracks, staff responsible for the operations of music products and complete beginners who are unfamiliar with related knowledge. We evaluate randomly sampled file, where each sample was evaluated by at least 5 raters.
@@ 6 Results
We begin by presenting our approach to refining captioning, which includes the capability for quality awareness, transitioning from text-level control to token-level control. Finally, we compare proposed model with previous works on both subjective and objective metrics.
6.1 Impact of music caption refinement
We conducted our ablation study on a subset of our training set, which includes ASM and FMA, totaling approximately 3,700 hours and 1.1 million clips. For evaluation, we utilized an out-of-domain set with 1,000 samples randomly selected from MTT (Law et al., 2009). Table  1 compares the model’s performance using different textual representations: sentences formed by merging original tags with commas ($\mathcal{D}_{o}$), generated captions ($\mathcal{D}_{g}$), and generated captions refined through filtering and fusion ($\mathcal{D}_{\text{merge}}$). During the filtering and fusion stage, 8.9% of the generated captions were filtered out, and 15.1% were fused with original tags using ChatGPT. Each model underwent training for 60,000 steps with a batch size of 64.
Table 1: Comparison of model performance training on different textual representations, evaluated by FAD, KL(Sigmoid based), IS and CLAP score.
@@ 6.1 Impact of music caption refinement
$\mathcal{D}_{o}$ 7.228 4.449 1.742 0.199 7.072 3.964 2.116 0.291
$\mathcal{D}_{g}$ 5.944 4.134 2.276 0.278 5.757 3.837 2.508 0.342
$\mathcal{D}_{\text{merge}}$ 5.867 4.132 2.294 0.284 5.637 3.804 2.625 0.350
From Table  1 we can also observe consistent trends: employing a captioner to transform audio annotations from sparse words into detailed sentences significantly improved the models’ generalization and diversity. This indicates that detailed annotations are essential for learning the relationship between the models and spectral features. Moreover, the filter and fusion stages led to enhancements across all metrics, highlighting the significance of precise, comprehensive annotations for generalization ability and control ability. We also found that compared to U-Net, the MDT architecture shows stable improvements in basic modeling metrics, making it a better backbone for music spectral modeling.
6.2 Quality awareness
This subsection aims to explore the effects and interactions of model control over quality tokens and quality text prefixes during the training phase, as well as their comparative effects across different models. In our previous MTT dataset of 1,000 test pairs, we filtered out pairs labeled with low quality or quality is poor to avoid confusion when acting with quality prefixes, resulting in a new subset of 519 entries, which we refer to as the MTT Filter Set (MTT-FS). Figure 3 shows the impact of different quality prefixes during inference when using quality as text prefix during training for U-Net and MDT based backbone. It is observed that U-Net, when infer with different quality prefixes, only shows minor changes in p-MOS scores and does not follow the threshold set for training. In contrast, MDT is better at learning quality information from prefixes, achieving p-MOS scores far above U-Net and the test set. Additionally, through decoupling of quality information from training set, we earn superior diversity and generalizability than training and inference without quality text prefix.
Given that quality tokens are specifically designed for the Transformer architecture, Figure 4 (left) demonstrates the controlled outcomes when different quality tokens are employed after integrating quantified quality as a token during training. Remarkably, using quality tokens alone provided more precise and accurate p-MOS score control.
For our ablation study, we compared the effects of solely using text prefixes against combining both approaches. As shown in Figure 4 (right), with given quantized control level increasing gradually, the model steadily improves p-MOS scores, which represent the quality of generation. Concurrently, FAD and KL also optimize progressively until a turning point occurs at level 4 while higher average p-MOS is reached than sorely using prefix. The turning point may be due to the scarcity of examples with quality level 5 in the dataset. Additionally, through combining two types of quality information injection, the refined decoupling and interaction allowed the model to more accurately perceive audio data quality features through training, leading to significantly reductions in FAD and KL compared to only using either of them. We also compare our approach with the traditional "Negative Prompt" strategy in Appendix A.3, highlighting our approach’s significant improvement in quality and reduction in FAD.
6.3 Compared with previous methods
We compared our proposed method with the following representative previous methods: AudioLDM 2 (Liu et al., 2023a) and MusicLDM (Chen et al., 2024b), which model music using spectral latent spaces, and MusicLM (Agostinelli et al., 2023), MeLoDy (Lam et al., 2024), and MusicGen Copet et al. (2024), which focus on modeling discrete representations. Additionally, we train AudioLDM (Liu et al., 2023c) on our own dataset with quality prefix for comparing between backbone designing and the use of quality token.
@@ 6.3 Compared with previous methods
Table 2: Objective evaluation results for music generation with diffusion based models and language model based models. We re-infer MusicLDM and AudioLDM 2 with label †.
Details Objective
Model Params Hours Fad $\downarrow$ Kl $\downarrow$ Is $\uparrow$ Clap $\uparrow$
MusicLM 1290M 280k 4.00 - - -
MusicGen 1.5B 20k 3.80 1.22 - -
Mousai 1042M 2.5k 7.50 1.59 - -
MeLoDy 726M 257.0k 5.41 - - -
MusicLDM† - 455.6 5.35 2.54 2.26 -
AudioLDM 2 - Full 712M 17.9k 3.13 1.20 - -
AudioLDM 2 - Music † 712M 10.8k 4.04 1.46 2.67 0.34
Ours (U-Net) 1.0B 12.5k 2.03 1.51 2.41 0.33
Ours (QA-MDT) 675M 12.5k 1.65 1.31 2.80 0.35
For AudioLDM 2, MusicLDM, we re-inferred based on their official links and checkpoints to compare additional metrics under the same environment. The results are shown in Table  2. Based on the experimental results, we achieved significant advantages in both subjective and objective metrics. Since the Kullback-Leibler divergence is used to measure the distance between audio samples, higher quality audio usually results in deviations from the original waveform of the dataset, thereby leading to lower performance on the KL metric.
@@
Table 3: Evaluation of model performances among different groups, rated for text relevance (Rel) and overall quality (Ovl), with higher scores indicating better performance. The groups included Production Operators (Po), Professional Music Producers (Pmp), and Video Editors (Ve).
Po Pmp Ve Beginners Average
Model Ovl Rel Ovl Rel Ovl Rel Ovl Rel Ovl Rel
Ground Truth 4.00 4.00 4.47 3.60 4.10 3.80 3.87 3.87 3.72 3.60
AudioLDM 2 2.03 2.42 3.03 3.61 3.21 3.71 3.85 3.85 3.03 3.40
Ours(U-Net) 2.80 3.34 3.46 4.08 3.40 3.96 3.88 3.96 3.38 3.84
Ours(QA-MDT) 3.27 3.77 3.69 4.19 3.54 3.94 4.23 4.00 3.68 3.97
@@
Based on subjective evaluation shown in Table 3, our proposed method significantly improves overall audio quality and text alignment, thanks to the label optimization scheme for large music datasets and the quality-aware training strategy. By analyzing the backgrounds of the evaluators and their corresponding results, we can also see that for beginners, the comparison between different systems is not sensitive, which is related to their lack of music background experience and knowledge. However, from the perspective of our method in product operators, video editors, and audio producers, our method offers considerable enhancements, underscoring its potential value to audio industry professionals.
@@ 7 Conclusion and discussion
In this study, we identify the challenges posed by large-scale uneven audio quality and unaligned textual annotations in the music generation domain, which hinder the development of diffusion-based TTM generation. By employing a novel quality decoupling learning approach based on p-MOS, along with masked diffusion transformer as the backbone for the diffusion process, we achieve enhanced generation quality and musicality for music generation.
Limitations.
Music, as a form of art, requires consideration of aesthetic qualities in generated melodies, in addition to quality. In this work, we did not explicitly optimize the generated melodies, nor did we explore the extrapolative performance of transformers on long-duration audio generation, which we will leave as future work.
Broader impact.
Except for the scaling ability, transformer-based generative models offer greater flexibility compared to U-Net, and the mask strategy facilitates tasks such as music inpainting and super-resolution spectral expansion. We hope our work paves the way for advancements in music generation, control, and other related applications.
@@ Appendix A Appendix
A.1 Training dataset.
Table A.1: Database statistics.
Database Duration Ave Length Clip Num Sample Rate Released
MTT 200h 30s 24K - 2009
FMA 900h 30s 10.9K 44.1kHz 2016
MSD 7333h 30s 880K 32kHz 2011
ASM 2777h 10s 1.0M mix 2017
Private 1375h 90s 55K mix -
Final 12.5kh 10s 45M 16kHz -
A.2 Ablation study for MDT structure
In image generation, for latent spaces with consistent aspect ratios, the patchify approach indicates that smaller square patch sizes ($2\times 2$ and $1\times 1$) achieve the best results [Peebles and Xie, 2023]. However, for our audio latent space, which has a size of $16\times 256$ with significant aspect ratio imbalance, the traditional square patch pattern may not be the best choice. Therefore, we chose $2\times 4$ as the fundamental patch size and conducted comparisons from three aspects: the decoder layer, patchify strategy, and in contrast to the traditional DiT.
Due to the potential inaccuracies inherent in tags, we evaluate our model on the Generated Captions without Filter and Fusion to assess its robustness against possibly inaccurate labels training with ASM and FMA, and test on MTT Test Set.
Decoder Layers
Table A.2: Performance comparison between numbers of decoder layers in the MTT Test Set.
Layers FAD $\downarrow$ KL $\downarrow$
6 5.937 3.912
8 5.901 3.912
10 5.958 3.724
To avoid introducing changes in the model’s parameter count, we fixed the total number of layers in the Encoder and Decoder at 28. As shown in the Table A.2, as the number of Decoder layers increases, the generality peaks at 8, and the sound quality gradually improves while diversity decreases. By conducting listening tests on a small subset of randomly sampled datasets, we found that the decrease in diversity leads to increasingly chaotic generated melodies, causing discomfort in the listening experience. Moreover, considering we have additional methods to control the quality of sound generation, to balance various metrics, we chose a layer count of 8 for subsequent experiments.
Pachify Strategy
In this section, we aim to explore the fundamental impact of different basic modeling units on spectral. From the experimental comparison of patch sizes $2\times 4$, $1\times 4$, and $2\times 2$ shown in Table A.3, it is evident that reducing the patch size consistently leads to performance improvements due to the more detailed spectral modeling. Considering the inherent spatial correlations in the Fourier transform within music spectrum, we analyze the results of experiments that apply spectral overlaps of 2 and 1 in the time and frequency domains separately, chich indicates that introducing overlap in the latent space does indeed contribute to improved results.
@@ A.2 Ablation study for MDT structure
A smaller patch sizes, such as $2\times 1$ or even the more extreme $1\times 1$, might offer improved performance. However, due to the significantly increased training and inference costs that accompany these sizes, we refrained from conducting further detailed experiments. Additionally, we explored the impact of these settings in comparison with DiT [Peebles and Xie, 2023, Chen et al., 2024a] architectures and discovered that the mask strategy, leveraging the spectral correlations in both the frequency and time domains, markedly enhances spectral modeling, which is evident both in terms of the convergence rate and the quality of the final results. However, based on subjective listening on a small test dataset, we find that although the overlap strategy can significantly improve the model’s objective indicators, it leads to a decline in both melodic generation and aesthetic aspects, which is crucial for music generation. Therefore, we choose $4\times 1$ as the base pachify strategy for subsequent experiments and the final model.
A.3 Comparing with using "negative prompt"
In this section, we aim to compare and analyze the traditional "negative prompt" method with our approach. In the implementation of CFG, we use
@@ A.3 Comparing with using "negative prompt"
where $y_{neg}$ is formulated as text embedding of "low quality."
@@
Based on the Table A.4, we found that any form of quality guidance improves the model’s generative performance. However, previous attempts to improve quality relied on the rare instances of "low quality" in the dataset. This necessitated the careful design of numerous negative prompts to avoid generating low-quality results. Furthermore, the text embedding of "low quality" might not be well disentangled during training, leading to an suboptimal results.

### v2

@@ 5 Experimental Setup
Datasets
Train datasets.
We used the following databases for our training: AudioSet Music Subset (ASM) (Gemmeke et al. 2017), MagnaTagTune (MTT) (Law et al. 2009), Million Song Dataset (MSD) (Bertin-Mahieux et al. 2011), Free Music Archive (FMA) (Defferrard et al. 2016), and an additional datasetWe use 55k music tracks from https://pixabay.com, which is a large scale copyright free dataset.. Each track in these databases was clipped to 10-second segments and sampled at 16kHz to ensure uniformity across the dataset. The final training set was developed through a process of caption refinement, as detailed in Section 4. Finally, we got our training set totaling 12.5k hours of diverse music data. The specific composition of these datasets is further elaborated in the Appendix.
Evaluation datasets.
For comparison with prior work, we evaluate our model on the widely used MusicCaps benchmark (Agostinelli et al. 2023) and the Song-Describer-Dataset (Manco et al. 2023). MusicCaps consists of 5.5K $10.24$-second clips that include high-quality music descriptions from ten musicians, and the Song-Describer-Dataset is made up of 706 licensed high quality music recordings.
Models and Hyperparameters
@@
Each 10.24-second audio clip, sampled at 16 kHz, is initially transformed into a $64\times 1024$ mel-spectrogram with mel-bins of 64, hop-length of 160 and window length of 1024. Subsequently, this spectrogram is compressed into a $16\times 128$ latent representation $\mathcal{M}_{spec}$ using a Variational Autoencoder (VAE) pretrained with AudioLDM 2 (Liu et al. 2023b) with series of quantization loss and adversarial loss. This latent representation is utilized for adding noise and training purposes. Finally, we employ a pretrained Hifi-GAN (Kong, Kim, and Bae 2020) to reconstruct the waveform from the generated mel-spectrogram.
Caption processing and conditioning.
@@
We utilize the LP-MusicCaps (Doh et al. 2023) caption model for ASM, FMA, and subsets of MTT and MSD that have weak or no captions. We use the official checkpoint from LAION-CLAP (Wu* et al. 2023)music_speech_audioset_epoch_15_esc_89.98.pt for text-to-text and text-to-audio similarity calculations. Based on small scale subjective experientment, thresholds are set at $\rho_{1}=\rho_{2}=0.1$ to ensure any generated text or original caption not aligned well with the corresponding waveform is filtered out. Additionally, after filtering, generated text that fall below a threshold of $\rho_{3}=0.25$ are merged with original tags with the prompt: Merge this music caption “generated caption” with the ground truth tags “original tags”. Do not add any imaginary elements and try to keep the modified caption refined and accurate. We use FLAN-T5-large (Peebles and Xie 2023) as text encoder for all models.
Diffusion backbone.
@@
We train our diffusion model with three backbones for comparison: U-Net (Ronneberger, Fischer, and Brox 2015) based at 1.0B parameters, code based on AudioLDM (Liu et al. 2023c); and our proposed Quality-aware Masked Diffusion Transformer (QA-MDT) based at 675M parameters with $N=20$ encoder layers and $M=8$ decoder layers. Flash attention (Dao et al. 2022) from the xFormers package (Lefaudeux et al. 2022) is utilized to improve training speed and memory usage for self-attention and cross-attention stages. We study the impact of the patch size and overlap size in the Appendix , and apply a patch size of $1\times 4$ without overlap for the training of our final model. We train on 10-second audio crops sampled at random from the full track, maintaining a total batch size of 64 on four NVIDIA A100 GPUs, with gradient clipping of 1.0, learning rate of 8e-5, 2000 steps of linear warmup without weight decay, and a condition drop of 0.1 during training. The final model was trained for a total of seven days (38.5k steps). During inference, we use Denoising Diffusion Implicit Models (DDIM) (Song, Meng, and Ermon 2020) with 200 steps and a guidance scale of 3.5, consistent with AudioLDM (Liu et al. 2023c). We begin by presenting our approach to refining captioning, which includes the capability for quality awareness, transitioning from text-level control to token-level control. Finally, we compare proposed model with previous works subjectively and objectively.
Evaluation Metrics
@@
We evaluate the proposed method using objective metrics, including the Fréchet Audio Distance (FAD) (Kilgour et al. 2018), Kullback-Leibler Divergence (KL), Inception Score (IS)All above metrics are computed using the audioldm_eval library (Liu et al. 2023c), ensuring standardized evaluation.. We also utilize pseudo-MOS scoring model (Ragano, Benetos, and Hines 2023) to estimate generation quality, with more accurate assessments derived from subjective metrics. For the human studies, we invite human raters to evaluate two aspects of the audio samples (i) overall quality (Ovl), and (ii) relevance to the text input (Rel). For the overall quality test, raters were asked to rate the perceptual quality of the provided samples with 1 to 5. For the text relevance test, raters were asked to rate the match between audio and text on a scale from 1 to 5. Our evaluators consist of people from various backgrounds, including professional music producers, video editors who need to use soundtracks, staff responsible for the operations of music products and complete beginners who are unfamiliar with related knowledge. We evaluate randomly sampled file, where each sample was evaluated by at least 5 raters.
@@ 6 Results
Quality Awareness
This subsection explores the effects and interactions of model control over quality tokens and quality text prefixes during the training phase, as well as their comparative effects across different models. In our previous MTT dataset of 1,000 test pairs, we filtered out pairs labeled with low quality or quality is poor to avoid confusion when applying quality prefixes, resulting in a new subset of 519 entries, which we refer to as the MTT Filter Set (MTT-FS). Figure 4 illustrates the impact of different quality prefixes during inference when quality is used as a text prefix during training for U-Net and MDT-based backbones. It was observed that U-Net, when inferred with different quality prefixes, showed only minor changes in p-MOS scores and did not adhere to the threshold set during training. In contrast, MDT demonstrated better learning of quality information from prefixes, achieving p-MOS scores significantly higher than those of U-Net and the test set. Additionally, by decoupling quality information from the training set, we achieved superior diversity and generalizability compared to training and inference without quality text prefixes. Given that quality tokens are specifically designed for the Transformer architecture, Figure 3 (left) shows the controlled outcomes when different quality tokens are used after integrating quantified quality as a token during training. Remarkably, using quality tokens alone provided more precise and accurate p-MOS score control. In our ablation study, we compared the effects of using only text prefixes against combining both approaches. As shown in Figure 3 (right), as the quantized control level gradually increased, the model steadily improved in p-MOS scores, which represent the quality of generation. Concurrently, FAD and KL also progressively optimized until a turning point at level 4, where a higher average p-MOS was achieved than when solely using prefixes. This turning point may be due to the scarcity of examples with quality level 5 in the dataset. Moreover, by combining two types of quality information injection, the refined decoupling and interaction allowed the model to more accurately perceive audio data quality features during training, leading to significant reductions in FAD and KL compared to using only one of the methods.
We also compare our approach with the traditional “negative prompt” strategy in Appendix, highlighting our approach’s significant improvement in quality and reduction in FAD.
Impact of Music Caption Refinement
Table 1: Comparison of model performance training on different textual representations, evaluated by FAD, IS and CLAP score.
@@
$\mathcal{D}_{o}$ 7.23 1.74 0.199 7.07 2.12 0.291
$\mathcal{D}_{g}$ 5.94 2.28 0.278 5.76 2.51 0.342
$\mathcal{D}_{\text{merge}}$ 5.87 2.29 0.284 5.64 2.63 0.350
@@
Table 2: Objective evaluation results for music generation with diffusion based models and language model based models. We re-infer AudioLDM 2 and MusicGen with label †.
Details MusicCaps Song Describer Dataset
Model Params Hours Fad $\downarrow$ Kl $\downarrow$ Is $\uparrow$ Clap $\uparrow$ Fad $\downarrow$ Kl $\downarrow$ Is $\uparrow$ Clap $\uparrow$
MusicLM 1290M 280k 4.00 - - - - - - -
MusicGen † 1.5B 20k 3.80 1.22 - 0.31 5.38 1.01 1.92 0.18
Mousai 1042M 2.5k 7.50 1.59 - 0.23 - - - -
Jen-1 746M 5.0k 2.0 1.29 - 0.33 - - - -
AudioLDM 2 - Full 712M 17.9k 3.13 1.20 - - - - - -
AudioLDM 2 - Music † 712M 10.8k 4.04 1.46 2.67 0.34 2.77 0.84 1.91 0.28
Ours (U-Net) 1.0B 12.5k 2.03 1.51 2.41 0.33 1.01 0.83 1.92 0.30
Ours (QA-MDT) 675M 12.5k 1.65 1.31 2.80 0.35 1.04 0.83 1.94 0.32
@@
Table 3: Evaluation of model performances among different groups, rated for text relevance (Rel) and overall quality (Ovl), with higher scores indicating better performance. The groups included Production Operators (Po), Professional Music Producers (Pmp), Video Editors (Ve) and Beginners(Bg)
Po Pmp Ve Bg
Model Ovl Rel Ovl Rel Ovl Rel Ovl Rel
Ground Truth 4.00 4.00 4.47 3.60 4.10 3.80 3.87 3.87
AudioLDM 2 2.03 2.42 3.03 3.61 3.21 3.71 3.85 3.85
MusicGen 2.83 3.54 2.63 2.92 3.41 3.00 4.33 3.83
Ours(U-Net) 2.80 3.34 3.46 4.08 3.40 3.96 3.88 3.96
Ours(QA-MDT) 3.27 3.77 3.69 4.19 3.54 3.94 4.23 4.00
We conducted our ablation study on a subset of our training set, which includes ASM and FMA, totaling approximately 3,700 hours and 1.1 million clips. For evaluation, we utilized an out-of-domain set with 1,000 samples randomly selected from MTT (Law et al. 2009). Table  1 compares the model’s performance using different textual representations: sentences formed by merging original tags with commas ($\mathcal{D}_{o}$), generated captions ($\mathcal{D}_{g}$), and generated captions refined through filtering and fusion ($\mathcal{D}_{\text{merge}}$). During the filtering and fusion stage, 8.9% of the generated captions were filtered out, and 15.1% were fused with original tags using ChatGPT. Each model underwent training for 60,000 steps with a batch size of 64.
From Table 1 we can also observe consistent trends: employing a captioner to transform audio annotations from sparse words into detailed sentences significantly improved the models’ generalization and diversity. This indicates that detailed annotations are essential for learning the relationship between the models and spectral features. Moreover, the filter and fusion stages led to enhancements across all metrics, highlighting the significance of precise, comprehensive annotations for generalization ability and control ability. We also found that compared to U-Net, the MDT architecture shows stable improvements in basic modeling metrics, making it a better backbone for music spectral modeling.
Compared with Previous Methods
We compared our proposed method with the following representative previous methods: AudioLDM 2 (Liu et al. 2023b), Mousai (Schneider et al. 2023) and Jen-1 (Li et al. 2023) which model music using spectral latent spaces, MusicLM (Agostinelli et al. 2023), and MusicGen (Copet et al. 2024), which focus on modeling discrete representations.
We re-inferred AudioLDM2-Music and MusicGen-1.5B using their official checkpoints to compare additional metrics under the same environment. The results are presented in Table 2. For Ours (U-Net), we inferred all text with the prefix “high quality”, while for Ours (QA-MDT), we used the same prefix along with a p-MOS quality token set to level 5. When calculating the CLAP score, we evaluated the generated music with original prompt, which did not include any quality prefix. The experimental results show significant advantages in both subjective and objective metrics for our models. Since KL divergence measures the distance between audio samples, higher quality audio often results in deviations from the original waveform of Musiccaps, which can lead to lower performance. Although Ours (U-Net) showed a slight FAD advantage on the Song-Describer-Dataset, this may be due to instabilities arising from the small dataset, and we further demonstrated the superiority of QA-MDT in subsequent subjective experiments. Additionally, since MusicGen was trained on non-vocal tracks, it may underperform on captions that include vocals.
@@
Based on subjective evaluation shown in Table 3, our proposed method significantly improves overall audio quality and text alignment, thanks to the label optimization for large music datasets and the quality-aware training strategy. By analyzing the backgrounds of the evaluators and their corresponding results, we can also see that for beginners, the comparison between different systems is not sensitive, which is related to their lack of music background experience and knowledge. However, from the perspective of our method in product operators, video editors, and audio producers, our method offers considerable enhancements, underscoring its potential value to audio industry professionals.
@@ 7 Conclusion and Discussion
In this study, we identify the challenges posed by large-scale uneven audio quality and unaligned textual annotations in the music generation domain, which hinder the development of diffusion-based TTM generation. By employing a novel quality awareness learning approach based on p-MOS, along with masked diffusion transformer as the backbone for the diffusion process, we achieve enhanced generation quality and musicality for music generation.
@@ Appendix A Appendix
Training Dataset
Table A.1: Database statistics. It should be noted that MSD is basically a package of commercial music, which is not a copyright-free content.
Database Duration (h) Clip Num Sample Rate Released
MTT 200 24K - 2009
FMA 900 10.9K 44.1kHz 2016
MSD 7333 880K 32kHz 2011
ASM 2777 1.0M mix 2017
Pixabay 1375 55K 44.1kHz -
Final 12.5K - 16kHz -
Patchify Strategy
In image generation, for latent spaces with consistent aspect ratios, the patchify approach indicates that smaller square patch sizes ($2\times 2$ and $1\times 1$) achieve the best results. However, for our audio latent space, which has a size of $16\times 256$ with significant aspect ratio imbalance, the traditional square patch pattern may not be the best choice. Therefore, we chose $2\times 4$ as the fundamental patch size and conducted comparisons from three aspects: the decoder layer, patchify strategy, and in contrast to the traditional DiT.
Due to the potential inaccuracies inherent in tags, we evaluate our model on the Generated Captions without Filter and Fusion to assess its robustness against possibly inaccurate labels training with ASM and FMA, and test on MTT Test Set. In this section, we aim to explore the fundamental impact of different basic modeling units on spectral. From the experimental comparison of patch sizes $2\times 4$, $1\times 4$, and $2\times 2$ shown in Table A.2, it is evident that reducing the patch size consistently leads to performance improvements due to the more detailed spectral modeling. Considering the inherent spatial correlations in the Fourier transform within music spectrum, we analyze the results of experiments that apply spectral overlaps of 2 and 1 in the time and frequency domains separately, which indicates that introducing overlap in the latent space does indeed contribute to improved results.
@@
A smaller patch sizes, such as $2\times 1$ or even the more extreme $1\times 1$, might offer improved performance. However, due to the significantly increased training and inference costs that accompany these sizes, we refrained from conducting further detailed experiments. Additionally, we explored the impact of these settings in comparison with DiT architectures and discovered that the mask strategy, leveraging the spectral correlations in both the frequency and time domains, markedly enhances spectral modeling, which is evident both in terms of the convergence rate and the quality of the final results. However, based on subjective listening on a small test dataset, we find that although the overlap strategy can significantly improve the model’s objective indicators, it leads to a decline in both melodic generation and aesthetic aspects, which is crucial for music generation. Therefore, we choose $1\times 4$ as the base patchify strategy for subsequent experiments and the final model.
Comparing with Using “Negative Prompt”
In this section, we aim to compare and analyze the traditional “negative prompt” method with our approach. In the implementation of CFG, we use
@@
where $y_{neg}$ is formulated as text embedding of ”low quality.”
@@
Based on the TableA.3, we found that any form of quality guidance improves the model’s generative performance. However, previous attempts to improve quality relied on the rare instances of ”low quality” in the dataset. This necessitated the careful design of numerous negative prompts to avoid generating low-quality results. Furthermore, the text embedding of ”low quality” might not be well disentangled during training, leading to an suboptimal results.